The tens digit of a 2 digit number is 2 larger than 3 times the units digit. If the sum of the digits is 10, find the number
Answer by jorel1380(3719)

m+n=10
3n+2=m
3n+2+n=10
4n=8
n=2
m=8
The number is 82..